name the vertex of the parabola with the equation y^2 - 8y + 18 = x
Answer by lwsshak3(11628)

name the vertex of the parabola with the equation
y^2-8y+18=x
y^2-8y=x-18
complete the square:
(y^2-8y+16)=x-18+16
(y-4)^2=(x-2)
This is an equation of a parabola that opens rightwards.
Its standard form of equation: (y-k)^2=4p(x-h), (h,k)=(x,y) coordinates of the vertex
For given parabola:
vertex: (2,4)